OpenStudy (jmoney132):

The radius of a cylinder is 5" and the lateral area is 70 pi sq. in. Find the length of the altitude.

OpenStudy (gnorris):

7 is the height or altitude

OpenStudy (gnorris):

Lateral Area = AL = 70*π Perimeter of Circle = PC = 2*π*R (R = Radius) Length of Altitude = L = AL / PC = 70*π / 2*π*R = 7 inches